# Generating Physically Stable and Buildable LEGO® Designs from Text

### v1

@@ 4 Method
Here, we introduce LegoGPT, a method for generating physically stable LEGO designs from text prompts. Leveraging LLMs’ ability to model sequences and understand text, we fine-tune a pre-trained LLM for the LEGO generation task (Section 4.1). To increase the stability and buildability of our designs, we use brick-by-brick rejection sampling and physics-aware rollback during inference (Section 4.2). Figure 3 illustrates an overview of our method.
@@ 4.1 Model Fine-tuning
Pre-trained LLMs excel at modeling sequences and understanding natural language, making them promising candidates for our task. We further fine-tune a pre-trained LLM on a custom instruction dataset containing text prompts and their corresponding LEGO structures from StableText2Lego.
Pre-trained Base Model.
We use LLaMA-3.2-1B-Instruct [11] as our base model. This model is fine-tuned to give coherent answers to instruction prompts, making it suitable for text-based LEGO design generation. As shown in Figure 5, the base model can generate LEGO-like designs through in-context learning, highlighting the promise of using pre-trained LLMs for our task. However, the generated designs often omit certain object parts and contain intersecting or disconnected bricks, making them physically unstable and unbuildable. To address these issues, we further fine-tune the pre-trained model using our dataset.
Instruction Fine-tuning Dataset.
For each stable design and its corresponding captions, we construct an instruction in the following format: “(user) Create a LEGO model of {caption}. (assistant) {lego-design}.”
To simplify training and reuse LLaMA’s tokenizer, we represent LEGO designs in plain text. But what format should we use? The standard LEGO design format LDraw [29] has two main drawbacks. First, it does not directly include brick dimensions, which are crucial for assessing the structure and validating brick placements. Second, it contains unnecessary information, such as brick orientation and scale. This information is redundant, as each axis-aligned brick has only two valid orientations.
@@
With our fine-tuned model LegoGPT $\theta$, we predict the bricks $b_{1},b_{2},...,b_{N}$ in an autoregressive manner:
@@ 4.2 Integrating Physical Stability
A LEGO structure is considered physically stable and buildable if it does not collapse when built on a baseplate. To this end, we assess physical structural stability using the stability analysis method [37]. Here, we briefly overview this method below. Figure 4(a) illustrates all possible forces exerted on a single brick. We extend the single brick model and derive the structural force model $\mathcal{F}$, which consists of a set of candidate forces (e.g., pulling, pressing, supporting, dragging, normal, etc.), as demonstrated in Figure 4(b). For a LEGO structure $B=[b_{1},b_{2},\dots,b_{N}]$, each brick $b_{i}$ has $M_{i}$ candidate forces $F_{i}^{j}\in\mathcal{F}_{i},j\in[1,M_{i}]$. A structure is stable if all bricks can reach static equilibrium, i.e.,
@@
$\begin{split}\operatorname*{arg\,min}_{\mathcal{F}}&\sum_{i}^{N}\Biggl{\{}|\sum_{j}^{M_{i}}F_{i}^{j}|+|\sum_{j}^{M_{i}}{\tau}_{i}^{j}|+\alpha\mathcal{D}_{i}^{\max}+\beta\sum\mathcal{D}_{i}\Biggl{\}},\end{split}$ (3)
@@
When to apply stability analysis?
Our model generates LEGO designs sequentially, one LEGO brick at a time. A straightforward approach to ensuring physical stability is to apply stability analysis to each step and resample a brick that would cause a collapse. However, this step-by-step validation, though efficient per check, could be time-consuming due to the large number of checks required. More importantly, many LEGO designs are unstable when partially constructed, yet become stable when fully assembled. Adding a stability check after each brick generation could overly constrain the model exploration space. Instead, we propose brick-by-brick rejection sampling combined with physics-aware rollback to balance stability and diversity.
Brick-by-Brick Rejection Sampling
To improve inference speed and avoid overly constraining the model generation, we relax our constraints during inference. First, when the model generates a LEGO brick and its position, the brick should be well-formatted (e.g., available in the inventory) and not lie outside the workspace. Second, we ensure that newly added bricks do not collide with the existing structure. Formally, for each generated brick $b_{t}$, we have $\mathcal{V}_{t}\cap\mathcal{V}_{i}=\varnothing,\forall i\in[1,t-1]$, where $\mathcal{V}_{i}$ denotes the voxels occupied by $b_{i}$. These heuristics allow us to efficiently generate well-formatted LEGO structures without explicitly considering complex physical stability. To integrate these heuristics, we use rejection sampling: if a brick violates the heuristics, we resample a new brick from the model. Due to the relaxed constraints, most bricks are valid, and rejection sampling does not significantly affect inference time.
Physics-Aware Rollback
To ensure that the final design $B=[b_{1},b_{2},\dots,b_{N}]$ is physically stable, we calculate the stability score $S$. If the resulting design is unstable, i.e., $s_{i}=0,i\in\mathcal{I}$, we roll back the design to the state before the first unstable brick was generated, i.e., $B^{\prime}=[b_{1},b_{2},\dots,b_{\min{\mathcal{I}}-1}]$. Here, $\mathcal{I}$ is the set of the indices of all the unstable bricks. We repeat this process iteratively until we reach a stable structure $B^{\prime}$, and continue generation from the partial structure $B^{\prime}$. Note that we can use the per-brick stability score to efficiently find the collapsing bricks and their corresponding indices in the sequence. We summarize our inference sampling in Algorithm 1.
@@
Output: LEGO design following the text prompt
1 $B\leftarrow\text{empty LEGO structure}$;
@@ 14                   $B\leftarrow[b_{1},\dots,b_{i-1}]$;
ALGORITHM 1 LegoGPT inference algorithm.
Table 1: Quantitative Analysis. We evaluate our method against several baselines on validity (no out-of-library, out-of-bounds, or colliding bricks), stability, and CLIP-based text similarity. Stability and CLIP score are computed over valid designs only. For LLaMA-Mesh [78], validity requires a well-formed OBJ file. Our method outperforms all baselines as well as the ablated setups on validity and stability using our proposed rejection sampling and rollback, while maintaining high text similarity.
Method % valid $\uparrow$ % stable $\uparrow$ mean brick stability $\uparrow$ min brick stability $\uparrow$ CLIP $\uparrow$
Pre-trained LLaMA (0-shot) 0.0% 0.0% N/A N/A N/A
In-context learning (5-shot) 2.4% 1.2% 0.675 0.479 0.284
LLaMA-Mesh [78] 94.8% 50.8% 0.894 0.499 0.311
LGM [68] 100% 25.2% 0.942 0.231 0.299
XCube [60] 100% 75.2% 0.964 0.686 0.322
Hunyuan3D-2 [86] 100% 75.2% 0.973 0.704 0.324
Ours w/o rejection sampling or rollback 37.2% 12.8% 0.956 0.325 0.329
Ours w/o rollback 100% 24.0% 0.947 0.228 0.322
Ours (LegoGPT) 100% 98.8% 0.996 0.915 0.324
Figure 5: Result gallery and baseline comparisons. Our method can generate high-quality, diverse, and novel LEGO designs aligned with the given text prompts. Black bricks are colliding. For LLaMA-Mesh [78], LGM [68], XCube [60], and Hunyuan3D-2 [86], an inset of the generated mesh before legolization is shown.
4.3 LEGO Texturing and Coloring
While our primary focus is on generating LEGO shapes, color and texture play a critical role in creative LEGO designs. Therefore, we propose a method that applies detailed UV textures or assigns uniform colors to individual bricks.
UV Texture Generation.
Given a structure $B$ and its corresponding mesh $\mathcal{M}$, we first identify the set of occluded bricks $B_{\text{occ}}$ that have all six faces covered by adjacent bricks, and remove $B_{\text{occ}}$ for efficiency. The remaining bricks $B_{\text{vis}}=B\setminus B_{\text{occ}}$ are merged into a single mesh $\mathcal{M}$ with cleaned overlapping vertices using ImportLDraw [72]. We generate a UV map $\text{UV}_{\mathcal{M}}$ by cube projection. The texture map $I_{\text{texture}}$ is then generated using FlashTex [9], a fast text-based mesh texturing approach:
@@ 4.3 LEGO Texturing and Coloring
Uniform Brick Color Assignment.
We can also assign each brick a uniform color from the LEGO color library. Given a structure $B$, we convert it to a voxel grid $\mathcal{V}$ and then to a UV-unwrapped mesh $\mathcal{M}_{\mathcal{V}}$. For every voxel $v\in\mathcal{V}$, let $f^{v}_{i},i=1,\dots,N_{v}$ be its visible faces where $0\leq N_{v}\leq 6$. Each face $f^{v}_{i}$ is split into two triangles and mapped to a UV region $\mathcal{S}^{v}_{i}$, creating a mesh $\mathcal{M}_{\mathcal{V}}$ with UV map $\text{UV}_{\mathcal{V}}$. We apply FlashTex [9] to generate a texture $I_{\text{texture}}$:
@@
where $\mathcal{C}(f^{v}_{i})=\frac{1}{|\mathcal{S}^{v}_{i}|}\sum_{(x,y)\in\mathcal{S}^{v}_{i}}I_{\text{texture}}(x,y)$ is the color of each visible face $f^{v}_{i}$, and $|\mathcal{S}^{v}_{i}|$ represents the number of pixels in region $\mathcal{S}^{v}_{i}$ in the UV map. For each brick $b_{t}$ and its constituent voxels $\mathcal{V}_{t}$, we compute the brick color $\mathcal{C}(b_{t})=\frac{1}{|\mathcal{V}_{t}|}\sum_{v\in\mathcal{V}_{t}}\mathcal{C}(v)$. Finally, we find the closest color in the color set. While UV texturing offers higher-fidelity details, uniform coloring allows us to use standard LEGO bricks.
Figure 6: Ablation study. Brick-by-brick rejection sampling and physics-informed rollback help to ensure that the generated LEGO design is both valid and stable. Black indicates colliding bricks.
Figure 7: Textured and Colored LEGO Generation. Our method can generate diverse textured (top two rows) and colored (bottom) LEGO models based on the same shape while using different appearance text prompts.
Figure 8: Automated Assembly. Demonstrations of automated assembly of generated LEGO structures using robots.
@@ 5.1 Implementation Details
Fine-tuning.
Our fine-tuning dataset contains 240k distinct prompts and 47k+ distinct LEGO designs. We use 90% of the data for training and 10% for evaluation. For efficiency, we include samples only up to 4096 tokens in length. Training details are provided in Appendix A.
Inference.
To evaluate our method, we generate one LEGO design for each of 250 prompts randomly selected from the validation dataset. The nonlinear optimization in Eqn. 3 is solved using Gurobi [17]. We set $F_{T}=0.98N$ with $\alpha=10^{-3}$ and $\beta=10^{-6}$. We allow up to 100 physics-aware rollbacks before accepting the LEGO design. The median number of required rollbacks is 2, and the median time to generate one LEGO design is 40.8 seconds.
5.2 LEGO Generation Results
Figure 5 shows a gallery of diverse, high-quality LEGO designs that closely follow the input prompts.
Baseline Comparisons.
As baselines, we use LLaMA-Mesh [78], LGM [68], XCube [60], and Hunyuan3D-2 [86] to generate a mesh from each prompt, and then convert the meshes to LEGO via legolization. Additionally, we compare our method with pre-trained models evaluated in both a zero-shot and few-shot manner. For few-shot evaluation, we provide the model with 5 examples of stable LEGO designs and their captions. We cannot run comparisons with Lennon et al. [30] as the code has not been released.
For each method, we compute the proportion of stable and valid structures among the generated designs. Additionally, for each valid structure, we compute its mean and minimum brick stability scores. To evaluate prompt alignment, we compute the CLIP score [59] between a rendered image of each valid structure and the text “A LEGO model of {prompt}”. As shown in Table 1, our method outperforms all baselines in these metrics. Figure 5 shows that our method generates LEGO structures of higher quality than the baselines.
Ablation Study.
We demonstrate the importance of rejection sampling and physics-aware rollback. As seen in Figure 6, rejection sampling eliminates invalid bricks, such as those with collisions, while rollback helps to ensure the final model is physically stable. The quantitative results in Table 1 show that our full method generates a higher proportion of valid and stable LEGO designs, while closely following the text prompts.
@@ 5.3 Extensions and Applications
Robotic Assembly of Generated LEGO.
We demonstrate automated assembly using a dual-robot-arm system in Figure 8. The robots use the manipulation policy [38] and the asynchronous multi-agent planner [23] to manipulate LEGO bricks and construct the structure. More details are included in Appendix B. Since the generated structures are physically stable, efficient and automated assembly can be performed.
Manual Assembly.
Our generated LEGO structures are physically valid and can be assembled by hand. See Appendix C for details.
Textured and Colored LEGO Model Generation.
Figure 7 shows both UV texturing and uniform coloring results of LEGO models, demonstrating our method’s ability to generate diverse styles while preserving the underlying geometry.

### v2

@@ 4 Method
Here, we introduce BrickGPT, a method for generating physically stable interconnecting toy brick assembly structures from text prompts. Leveraging LLMs’ ability to model sequences and understand text, we fine-tune a pre-trained LLM for the brick structure generation task (Section 4.1). To increase the stability and buildability of our designs, we use brick-by-brick rejection sampling and physics-aware rollback during inference (Section 4.2). Figure 3 illustrates an overview of our method.
@@ 4.1 Model Fine-tuning
Pre-trained LLMs excel at modeling sequences and understanding natural language, making them promising candidates for our task. We further fine-tune a pre-trained LLM on a custom instruction dataset containing text prompts and their corresponding brick structures from StableText2Brick.
Pre-trained Base Model.   We use LLaMA-3.2-1B-Instruct [11] as our base model. This model is fine-tuned to give coherent answers to instruction prompts, making it suitable for text-based brick structure generation. As shown in Figure 5, the base model can generate brick structures through in-context learning, highlighting the promise of using pre-trained LLMs for our task. However, the generated structures fail to follow the input prompt, and they contain intersecting or disconnected bricks, making them physically unstable and unbuildable. To address these issues, we further fine-tune the pre-trained model using our StableText2Brick.
Instruction Fine-tuning Dataset.   For each stable structure and its corresponding captions, we construct an instruction in the following format: “(user) Create a LEGO model of {caption}. (assistant) {brick-structure}.”
To simplify training and reuse LLaMA’s tokenizer, we represent brick structures in plain text. But what format should we use? The standard format LDraw [30] has two main drawbacks. First, it does not directly include brick dimensions, which are crucial for assessing the structure and validating brick placements. Second, it contains unnecessary information, such as brick orientation and scale. This information is redundant, as each axis-aligned brick has only two valid orientations.
@@
With our fine-tuned BrickGPT model $\theta$, we predict the bricks $b_{1},b_{2},...,b_{N}$ in an autoregressive manner:
@@ 4.2 Integrating Physical Stability
A brick structure is considered physically stable and buildable if it does not collapse when built on a baseplate. To this end, we assess physical structural stability using the stability analysis method [38]. We briefly overview this method below. Figure 4(a) illustrates all possible forces exerted on a single brick. We extend the single brick model and derive the structural force model $\mathcal{F}$, which consists of a set of candidate forces (e.g., pulling, pressing, supporting, dragging, normal, etc.), as shown in Figure 4(b). For a brick structure $B=[b_{1},b_{2},\dots,b_{N}]$, each brick $b_{i}$ has $M_{i}$ candidate forces $F_{i}^{j}\in\mathcal{F}_{i},j\in[1,M_{i}]$. A structure is stable if all bricks can reach static equilibrium, i.e.,
@@
$\begin{split}\operatorname*{arg\,min}_{\mathcal{F}}&\sum_{i}^{N}\Biggl{\{}\bigg{|}\sum_{j}^{M_{i}}F_{i}^{j}\bigg{|}+\bigg{|}\sum_{j}^{M_{i}}{\tau}_{i}^{j}\bigg{|}+\alpha\mathcal{D}_{i}^{\max}+\beta\sum\mathcal{D}_{i}\Biggl{\}},\end{split}$ (3)
@@
When to apply stability analysis?   Our model generates structures sequentially, one brick at a time. A straightforward approach to ensuring physical stability is to apply stability analysis to each step and resample a brick that would cause a collapse. However, this step-by-step validation, though efficient per check, could be time-consuming due to the large number of checks required. More importantly, many structures are unstable when partially constructed, yet become stable when fully assembled. Adding a stability check after each brick generation could overly constrain the model exploration space. Instead, we propose brick-by-brick rejection sampling combined with physics-aware rollback to balance stability and diversity.
Brick-by-Brick Rejection Sampling.   To improve inference speed and avoid overly constraining the model generation, we relax our constraints during inference. First, when the model generates a brick and its position, the brick should be well-formatted (e.g., available in the inventory) and not lie outside the workspace. Second, we ensure that newly added bricks do not collide with the existing structure. Formally, for each generated brick $b_{t}$, we have $\mathcal{V}_{t}\cap\mathcal{V}_{i}=\varnothing,\forall i\in[1,t-1]$, where $\mathcal{V}_{i}$ denotes the voxels occupied by $b_{i}$. These heuristics allow us to efficiently generate well-formatted brick structures without explicitly considering complex physical stability. To integrate these heuristics, we use rejection sampling: if a brick violates the heuristics, we resample a new brick from the model. Due to the relaxed constraints, most bricks are valid, and rejection sampling does not significantly affect inference time.
Physics-Aware Rollback.   To ensure that the final design $B=[b_{1},b_{2},\dots,b_{N}]$ is physically stable, we calculate the stability score $S$. If the resulting design is unstable, i.e., $s_{i}=0,i\in\mathcal{I}$, we roll back the design to the state before the first unstable brick was generated, i.e., $B^{\prime}=[b_{1},b_{2},\dots,b_{\min{\mathcal{I}}-1}]$. Here, $\mathcal{I}$ is the set of the indices of all the unstable bricks. We repeat this process iteratively until we reach a stable structure $B^{\prime}$, and continue generation from the partial structure $B^{\prime}$. Note that we can use the per-brick stability score to efficiently find the collapsing bricks and their corresponding indices in the sequence. We summarize our inference sampling in Algorithm 1.
@@
Output: Brick structure following the text prompt
1 $B\leftarrow\text{empty brick structure}$;
@@ 14                   $B\leftarrow[b_{1},\dots,b_{i-1}]$;
ALGORITHM 1 BrickGPT inference algorithm.
Table 1: Quantitative Analysis. We evaluate our method against several baselines on validity (no out-of-library, out-of-bounds, or colliding bricks), stability, CLIP-based text similarity, and DINOv2-based image similarity. Stability, CLIP, and DINO are computed over valid structures only. For LLaMA-Mesh [81], validity requires a well-formed OBJ file. Our method outperforms all baselines as well as the ablated setups on validity and stability using our proposed rejection sampling and rollback, while maintaining high text similarity.
Method % valid % stable mean brick stability min brick stability CLIP DINO
Pre-trained LLaMA (0-shot) 0.0% 0.0% N/A N/A N/A N/A
In-context learning (5-shot) 2.4% 1.2% 0.675 0.479 0.284 0.814
LLaMA-Mesh [81] 94.8% 50.8% 0.894 0.499 0.317 0.851
LGM [70] 100% 25.2% 0.942 0.231 0.300 0.851
XCube [62] 100% 75.2% 0.964 0.686 0.322 0.859
Hunyuan3D-2 [89] 100% 75.2% 0.973 0.704 0.324 0.868
Ours w/o rejection sampling or rollback 37.2% 12.8% 0.956 0.325 0.329 0.888
Ours w/o rollback 100% 24.0% 0.947 0.228 0.322 0.882
Ours (BrickGPT) 100% 98.8% 0.996 0.915 0.324 0.880
Figure 5: Result gallery and baseline comparisons. Our method generates high-quality, diverse, and novel brick structures aligned with the given text prompts. Black bricks are colliding. For LLaMA-Mesh [81], LGM [70], XCube [62], and Hunyuan3D-2 [89], an inset of the generated mesh is shown in the top-left corner.
4.3 Brick Texturing and Coloring
While we primarily focus on generating the shape of a brick structure, color and texture play a critical role in creative designs. Therefore, we propose a method that applies detailed UV textures or assigns uniform colors to individual bricks.
UV Texture Generation.   Given a structure $B$ and its corresponding mesh $\mathcal{M}$, we first identify the set of occluded bricks $B_{\text{occ}}$ that have all six faces covered by adjacent bricks, and remove $B_{\text{occ}}$ for efficiency. The remaining bricks $B_{\text{vis}}=B\setminus B_{\text{occ}}$ are merged into a single mesh $\mathcal{M}$ with cleaned overlapping vertices using ImportLDraw [74]. We generate a UV map $\text{UV}_{\mathcal{M}}$ by cube projection. The texture map $I_{\text{texture}}$ is then generated using FlashTex [8], a fast text-based mesh texturing approach:
@@ 4.3 Brick Texturing and Coloring
Uniform Brick Color Assignment.   We can also assign each brick a uniform color from a standard color library [30]. Given a structure $B$, we convert it to a voxel grid $\mathcal{V}$ and then to a UV-unwrapped mesh $\mathcal{M}_{\mathcal{V}}$. For every voxel $v\in\mathcal{V}$, let $f^{v}_{i},i=1,\dots,N_{v}$ be its visible faces where $0\leq N_{v}\leq 6$. Each face $f^{v}_{i}$ is split into two triangles and mapped to a UV region $\mathcal{S}^{v}_{i}$, creating a mesh $\mathcal{M}_{\mathcal{V}}$ with UV map $\text{UV}_{\mathcal{V}}$. We apply FlashTex [8] to generate a texture $I_{\text{texture}}$:
@@
where $\mathcal{C}(f^{v}_{i})=\frac{1}{|\mathcal{S}^{v}_{i}|}\sum_{(x,y)\in\mathcal{S}^{v}_{i}}I_{\text{texture}}(x,y)$ is the color of each visible face $f^{v}_{i}$, and $|\mathcal{S}^{v}_{i}|$ represents the number of pixels in region $\mathcal{S}^{v}_{i}$ in the UV map. For each brick $b_{t}$ and its constituent voxels $\mathcal{V}_{t}$, we compute the brick color $\mathcal{C}(b_{t})=\frac{1}{|\mathcal{V}_{t}|}\sum_{v\in\mathcal{V}_{t}}\mathcal{C}(v)$. Finally, we find the closest color in the color set. While UV texturing offers higher-fidelity details, uniform coloring allows us to use standard toy bricks.
Figure 6: Ablation study. Brick-by-brick rejection sampling and physics-informed rollback help to ensure that the generated structure is both valid and stable. Black indicates colliding bricks.
Figure 7: Brick Texture and Color Generation. Our method can generate diverse textured (top two rows) and colored (bottom) brick structures based on the same shape while using different appearance text prompts.
Figure 8: Automated Assembly. We demonstrate robotic assembly of generated structures using LEGO bricks.
@@ 5.1 Implementation Details
Fine-tuning.   Our fine-tuning dataset contains 240k distinct prompts and 47k+ distinct brick structures. We use 90% of the data for training and 10% for evaluation. For efficiency, we include samples only up to 4096 tokens in length. Training details are provided in Appendix A.
Inference.   To evaluate our method, we generate one brick structure for each of 250 prompts randomly selected from the validation dataset. The nonlinear optimization in Eqn. 3 is solved using Gurobi [18]. We set $F_{T}=0.98N$ with $\alpha=10^{-3}$ and $\beta=10^{-6}$. We allow up to 100 physics-aware rollbacks before accepting the brick structure. The median number of required rollbacks is 2, and the median time to generate one structure is 40.8 seconds.
5.2 Brick Structure Generation Results
Figure 5 shows a gallery of diverse, high-quality brick structures that closely follow the input prompts.
Baseline Comparisons.   As baselines, we use LLaMA-Mesh [81], LGM [70], XCube [62], and Hunyuan3D-2 [89] to generate a mesh from each prompt, then convert the meshes to brick structures with our delete-and-rebuild algorithm. Additionally, we compare our method with pre-trained LLaMA models evaluated in both a zero-shot and few-shot manner. For few-shot evaluation, we provide the model with 5 examples of stable brick structures and their captions.
For each method, we compute the proportion of stable and valid structures among the generated designs. Additionally, for each valid structure, we compute its mean and minimum brick stability scores. To evaluate prompt alignment, we compute the CLIP score [61] between a rendered image of each valid structure and the text “A LEGO model of {prompt}”. We also calculate the alignment between rendered images of the generated structure and the ground-truth structure for the same prompt, as measured by the cosine similarity between DINOv2 features [59]. As shown in Table 1, our method outperforms all baselines in these metrics. Figure 5 shows that our method generates brick structures of higher quality than the baselines.
Ablation Study.   We demonstrate the importance of rejection sampling and physics-aware rollback. As seen in Figure 6, rejection sampling eliminates invalid bricks, such as those with collisions, while rollback helps to ensure the final assembly structure is physically stable. The quantitative results in Table 1 show that our full method generates a higher proportion of valid and stable brick structures, while closely following the text prompts.
@@ 5.3 Extensions and Applications
Robotic Assembly of Generated Structures.   We demonstrate automated assembly using a dual-robot-arm system in Figure 8. The robots use the manipulation policy [39] and the asynchronous multi-agent planner [24] to manipulate toy bricks and construct the structure. More details are included in Appendix B. Since the generated structures are physically stable, efficient and automated assembly can be performed.
Manual Assembly.   Our generated brick structures are physically valid and can be assembled by hand. See Appendix C for details.
Brick Texture and Color Generation.   Figure 7 shows both UV texturing and uniform coloring results of brick structures, demonstrating our method’s ability to generate diverse styles while preserving the underlying geometry.